# The Role of Graph Topology in the Performance of Biomedical Knowledge Graph Completion Models

### v1

@@ 5.1 Topological Properties of Biomedical KGs
We first investigate how prevalent the different topological patterns, as described in Section 3, are in the range of public biomedical KGs we consider. Analysing this, we find that many-to-many is the predominant edge cardinality in all considered KGs, with only scarce occurrence of one-to-one, one-to-many, and many-to-one triples (Figure 3). This is in line with the majority of triples having an head out-degree and tail in-degree larger than one (Figures 4 and A.1).
@@
Table 1: Occurrence of edge topological patterns as fraction of total triples in the datasets. Note that, since PrimeKG was pre-processed to remove reverse edges as described in Section 4, it does not contain any symmetric triples.
Graph Sym Inf Inv Comp
Hetionet $0.002$ $0.124$ $0.001$ $0.693$
OpenBioLink $0.317$ $0.372$ $0.359$ $0.840$
[…]
PharmKG $0.197$ $0.124$ $0.059$ $0.651$
[…]
FB15k-237 $0.113$ $0.161$ $0.217$ $0.645$
@@ 5.2 Effect of Topological Properties on Model Accuracy
Overall, we observe a significant variance in mean reciprocal rank (MRR) achieved between the different KGs as well as between KGE models (Figure 5). In an attempt to understand the root cause of this variance, we focus our analysis on the differences in edge cardinalities and topological patterns described in Section 5.1. When only considering the average occurrence of these topological properties per dataset, the data supports no conclusion about the effect on model accuracy of either edge cardinalities (Figure 1(a)) or edge topological patterns (Figure 1(b)). Previous work (e.g., Teneva & Hruschka (2023)) went one step further, classifying relation types based on the predominant edge topological pattern and cardinality and evaluating the model accuracy per relation type. However, this also does not result in a clear connection between topological properties and KGE model accuracy (Figure C.2). We hypothesize that this could be due the confounding effects of covariates, such as node degrees and different topological patterns, that are too difficult to disentangle, as these properties are often not homogenous within a given relation type. Therefore, unlike prior work, we dissect the link between topological properties and model accuracy at the level of individual triples to allow for a finer-grained analysis with an improved statistical power and a reduced impact of covariates.
@@
When comparing the edge cardinality of individual triples to the MRR achieved by the models on these triples, the observed effect strongly differs between the investigated KGs (Figure 6). This result suggests that the exact entity degrees, together with other topological properties, might be better suited to explain the MRR than a binary one/many cardinality classification.
@@
Indeed, for a given triple, we observe a strong positive correlation between model accuracy and the in-degree of the tail entity as well as a strong negative correlation between model accuracy and the out-degree of the head entity (Figure C.3). This is observed even more clearly when considering the degrees of same relation type (Figures 7 and C.4; see Figures 4 and A.1 for the distribution of degrees in the different KGs). In fact, a high in-degree of the tail node in a tail prediction task has been linked to a higher score (Bonner et al., 2022b), therefore increasing the likelihood of the model predicting it, as confirmed in Figure C.5. On the other hand, a high out-degree of same relation of the head node in a query $(h,r,?)$, implies multiple correct tail entities. Due to the incompleteness of the KG, some of these correct triples might not be present in the graph and thus cannot be filtered out during inference, making the task of predicting the specific entity that is expected by the model more challenging.
@@
Consequently, in an attempt to adjust for entity degree, we investigate the effect of edge topological patterns within sets of triples binned by head and tail degrees. Overall, we find that the impact of edge topological patterns on model performance is more relevant if the degrees of head and tail entities are small. When this is the case, compositions are beneficial to MRR across all datasets and models (Figure 8 and Figure C.6). Interestingly, this is true even for DistMult that can’t explicitly model compositional patterns.
Figure 8: The effect of having compositions on MRR, triples grouped by their head and tail degrees of same relation.
When analysing the effect of the patterns is symmetric, has inference and has inverse, we need to distinguish whether the counterpart of the tested edge was present in the training split. If so, the prediction typically becomes much easier due to the availability of additional information seen during training (Toutanova & Chen, 2015), as shown in Figure C.7. An exception to this rule is the prediction of symmetric triples with the scoring function TransE: since TransE by design cannot model symmetric relations (Table A.2), it performs poorly on symmetric triples independently of the counterpart being present in the training data or not. This factor is likely contributing to the subpar accuracy of TransE on OpenBioLink and PharmKG, which have a large fraction of symmetric triples (Figures 5 and 1). On the other hand, our experiments confirm an above average performance of DistMult (which models all relations as symmetric) on symmetric triples, see Figure C.7. In the more challenging case, where the counterpart edge is not present in the training data, symmetry shows a detrimental effect on MRR whereas having inverse or inference has only little impact on accuracy (Figure C.7).
@@ 5.3 Predicting Specific Relation Types
To apply the analysis conducted in the previous sections to real-world tasks, we focus on the relations between pairs of entity types that are of particular interest to practitioners (Gene-Gene, Drug-Gene and Drug-Disease). We investigate how they are represented in different biomedical KGs, and how these differences affect the predictive performance of KGE models. Statistics for the interactions of the considered entity types are given in Figure C.8, while MRR is plotted in Figure 9.
One basic way of characterising the performance of a KGE model is to consider if it can rank entities of the correct type above those of other types when predicting a specific relation (for instance, gene entities when predicting drug-gene interactions) – a task we refer to as demixing. As displayed in Figure C.9, the trained models tend to consistently predict entities of the correct type, albeit with different levels of accuracy. Interestingly, DistMult often shows worse demixing capabilities than other scoring functions. As a consequence of the generally strong demixing performance, the size of the potential set of tails for a given interaction needs to be taken into consideration when comparing MRR across different datasets and relation types (the smaller the set of candidates, the higher the expected MRR achieved by random ranking).
Gene-Gene. As shown in Figure 8(a), these interactions are characterized by a large number of potential tails and a large head out-degree of same relation (with the exception of PharmKG for the latter), making predictions hard. However, a high MRR is observed for PharmKG and OpenBioLink: this is explained by the fact that, for these datasets only (Figure 8(b)), most triples are symmetric (and a significant portion also have inverse/inference edges), with the counterpart likely to have been seen during training (Section 5.2). This also explains the relative ordering of scoring functions for OpenBioLink and PharmKG, with DistMult being the best and TransE the worst.
Drug-Gene. In Hetionet, PrimeKG and OpenBioLink these, interactions are easier to predict for the models than gene-gene interactions (Figure 9). This is likely due to an overall smaller number of potential tails seen during training and, in the case of PrimeKG, a markedly smaller head out-degree. PharmKG and PharMeBINet do not satisfy either of these properties, and indeed we observe no improvements in MRR. The strong predictive power observed for OpenBioLink can be linked to the presence of inverse edges for the vast majority of these triples.
@@ 5.4 Case Study: Effect of Additional Training Data
We consider eight relation types where we can find a significant number of common triples between the two datasets (statistics are given in Table C.1). We extract 10% of the shared triples of the considered relation type as test set and use all other edges in each graph for training. In the case of Hetionet, in addition to the embedding size selected in Section 5.2 maximizing memory utilization, we also train models with the same embedding size chosen for PharMeBINet, which is generally strictly smaller (Hetionet_max and Hetionet_same in Figure 10; hyperparameters in Table B.1). To further ensure a fair comparison, at test time we restrict predictions to a custom set of candidate tails that is the same for the two datasets, namely the set of entities appearing as tails in shared triples of the given relation type.
@@
As shown in Figures 10 and C.10, across all tested relation types there is no indication that the KGE models are able to benefit from the additional data seen when training on the larger PharMeBINet graph. On the contrary, models trained on Hetionet consistently perform better, even when using the same embedding size. Interestingly, the gap in MRR varies strongly across relation types and also across scoring functions, with DistMult showing generally little difference when comparing models trained with the same embedding size, while distance-based scoring functions experience worse degradation. From Table C.1 we notice that the relation types where the MRR gap is more significant tend to have a larger overall head out-degree (and more relation types coming out from head nodes) for triples in PharMeBINet, compared to Hetionet. We hypothesise that the fact that these nodes are used as head entities for more relation types and triples, many of which are likely not relevant for the specific task at hand, negatively impacts the quality of embeddings. On the other hand, Compound-downregulates-Gene and Compound-causes-Side Effect exemplify relation types where PharMeBINet contains far more triples in addition to the ones in Hetionet. Even in this case, where we could expect the additional training data in the larger dataset to be strictly relevant to the prediction task, all scoring functions show markedly degraded performance when trained on PharMeBINet.
@@ 6 Conclusions
This paper analyses the topological properties of widely used biomedical KGs and compares the corresponding predictive performance of different KGE models, focussing in particular on link-prediction tasks relevant to practitioners. Deviating from previous studies, we justify the need to look at properties of individual edges when trying to explain results, as pooling at the relation level introduces too much noise to detect any patterns. By going beyond the coarse one/many binary classification typically used for edge cardinality, we find that considering the actual degrees of head and tail nodes gives a stronger predictor explaining model accuracy. Interestingly, similar interpretations apply to both biomedical KGs and the general-domain KG. Edge topological patterns also impact predictive accuracy, especially when entity degrees are small. For such patterns, we observe an improved accuracy when the counterpart edge (e.g., the reverse edge for symmetric triples) has been seen during training. This in turn raises the problem of inconsistent model validation schemes and topological property definitions when comparing results of previous studies on similar datasets. We address this problem by releasing all predictions from our experiments, together with a new toolbox for KG topological analysis. Finally, by performing a case study comparing predictions on common sets of edges shared by different KGs, we show that training on larger graphs, encoding more biomedical data, can unintuitively harm predictive performance. This evidence should encourage a wider discussion on the guiding principles to adopt when constructing the training KG for biomedical tasks – a crucial problem for real-world practitioners, but scarcely investigated in the literature.
@@ Appendix A Dataset and Model Properties
Table A.2: Scoring functions and their ability to model four fundamental relation properties: S = Symmetry; INF = Inference; INV = Inversion; C = Composition. For RotatE we assume $d$ even and denote by $\mathbb{C}^{\frac{d}{2}}$ the vector space $\mathbb{R}^{d}=(\mathbb{R}\oplus i\mathbb{R})^{\frac{d}{2}}$ with the structure of $\mathbb{R}$-algebra induced by the product of complex numbers. $\circ$ denotes the Hadamard product; $p\in\left\{1,2\right\}$.
@@
TransE $-\left\lVert{\bm{h}}+{\bm{r}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{r}},{\bm{t}}\in\mathbb{R}^{d}$
RotatE $-\left\lVert{\bm{h}}\circ e^{i{\bm{r}}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{t}}\in\mathbb{C}^{\frac{d}{2}},{\bm{r}}\in\mathbb{R}^{\frac{d}{2}}$
@@
TripleRE $-\left\lVert{\bm{h}}\circ\bm{r_{h}}+\bm{r_{m}}-{\bm{t}}\circ\bm{r_{t}}\right\rVert$ ${\bm{h}},\bm{r_{h}},\bm{r_{m}},{\bm{t}},\bm{r_{t}}\in\mathbb{R}^{d}$
@@ Appendix B Details on Experimental Setup and Hyperparameter Selection
[…]
@@
Hetionet DistMult $2048$ $0.0003$ $16$
[…]
Hetionet_same DistMult $300$ $0.0003$ $16$
[…]
OpenBioLink DistMult $768$ $0.0003$ $16$
[…]
PharMeBINet DistMult $300$ $0.003$ $16$
[…]
PharmKG DistMult $768$ $0.003$ $16$
[…]
PrimeKG DistMult $1024$ $0.0003$ $16$
[…]
FB15k-237 DistMult $4096$ $0.001$ $16$
[…]

### v2

@@ 5.1 Topological Properties of Biomedical KGs
We first investigate how prevalent the different topological patterns are in the range of public biomedical KGs we consider. Analysing this, we find that many-to-many is the predominant edge cardinality in all considered KGs, with only scarce occurrence of one-to-one, one-to-many, and many-to-one triples (Figure 4). This is in line with the majority of triples having an head out-degree and tail in-degree larger than one (Figure 5 and Figure A.1).
@@
Table 1: The fraction of edges with a given topological pattern (symmetry, inference, inverse, composition, or none of the four) in the datasets. As a single edge can satisfy multiple patterns, the sum of the frequencies in each row can exceed 1. Note that, since PrimeKG was pre-processed to remove reverse edges as described in Experimental Setup, it does not contain any symmetric triples.
Graph Sym Inf Inv Comp None
Hetionet $0.002$ $0.124$ $0.001$ $0.693$ $0.298$
OpenBioLink $0.317$ $0.372$ $0.359$ $0.840$ $0.12$
PharMeBINet 2.42e42.4210-4 $0.052$ $0.002$ $0.598$ $0.396$
PharmKG $0.197$ $0.124$ $0.059$ $0.651$ $0.279$
PrimeKG 0 2.08e42.0810-4 0 $0.807$ $0.193$
FB15k-237 $0.113$ $0.161$ $0.217$ $0.645$ $0.287$
@@ 5.2 Effect of Topological Properties on Model Accuracy
Overall, we observe a significant variance in mean reciprocal rank (MRR) achieved between the different KGs as well as between KGE models (Figure 6; see also Figure C.1 for Hits@1 and Hits@10, where we observe similar variability and consistent trends in the results). In an attempt to understand the root cause of this variance, we focus our analysis on the differences in edge cardinalities and topological patterns. When only considering the average occurrence of these topological properties per dataset, the data supports no conclusion about the effect on model accuracy of either edge cardinalities (Figure 3(a)) or edge topological patterns (Figure 3(b)). Previous work (e.g., (Teneva & Hruschka, 2023)) went one step further, classifying relation types based on the predominant edge topological pattern and cardinality and evaluating the model accuracy per relation type. However, this also does not result in a clear connection between topological properties and KGE model accuracy (Figure C.4). We hypothesize that this could be due to the confounding effects of covariates, such as node degrees and different topological patterns, that are too difficult to disentangle, as these properties are often not homogenous within a given relation type. Therefore, unlike prior work, we dissect the link between topological properties and model accuracy at the level of individual triples to allow for a finer-grained analysis with an improved statistical power and a reduced impact of covariates.
@@
When comparing the edge cardinality of individual triples to the MRR achieved by the models on these triples, the observed effect strongly differs between the investigated KGs (Figure 7, and similarly for Hits@10 in Figure C.2). This result suggests that the exact entity degrees, together with other topological properties, might be better suited to explain the model accuracy than a binary one/many cardinality classification.
@@
Indeed, for a given triple, we observe a strong positive correlation between model accuracy and the in-degree of the tail entity as well as a strong negative correlation between model accuracy and the out-degree of the head entity (Figure C.5). This is observed even more clearly when considering the degrees of same relation type (Figure 8 and Figures C.6 and C.7; the same patterns are observed for Hits@10, as exemplified in Figure C.8). In fact, a high in-degree of the tail node in a tail prediction task has been linked to a higher score (Bonner et al., 2022b), therefore increasing the likelihood of the model predicting it, as confirmed in Figure C.9. On the other hand, a high out-degree of same relation type of the head node in a query $(h,r,?)$, implies multiple correct tail entities. Due to the incompleteness of the KG, some of these correct triples might not be present in the graph and thus cannot be filtered out during inference, making the task of predicting the specific entity that is expected by the model more challenging.
@@
Consequently, in an attempt to adjust for entity degree, we investigate the effect of edge topological patterns within sets of triples binned by head and tail degrees. Overall, we find that the impact of edge topological patterns on model performance is more relevant if the degrees of head and tail entities are small. When this is the case, compositions are beneficial to model accuracy across all datasets and models (see Figure 9 and Figures C.10 and C.11 for MRR; the same holds for Hits@10, e.g., Figure C.12). Interestingly, this is true even for DistMult that can’t explicitly model compositional patterns.
Figure 9: The effect of having compositions on MRR, triples grouped by their head and tail degrees of same relation type.
When analysing the effect of the patterns is symmetric, has inference and has inverse, we need to distinguish whether the counterpart of the tested edge was present in the training split. If so, the prediction typically becomes much easier due to the availability of additional information seen during training (Toutanova & Chen, 2015), as shown in Figure C.13. An exception to this rule is the prediction of symmetric triples with the scoring function TransE: since TransE by design cannot model symmetric relations (Table A.2), it performs poorly on symmetric triples independently of the counterpart being present in the training data or not. This factor is likely contributing to the subpar accuracy of TransE on OpenBioLink and PharmKG, which have a large fraction of symmetric triples (Figures 6 and 1). On the other hand, our experiments confirm an above average performance of DistMult (which models all relations as symmetric) on symmetric triples, see Figure C.13. In the more challenging case, where the counterpart edge is not present in the training data, symmetry shows a detrimental effect on model accuracy, whereas having inverse or inference has only little impact on accuracy (Figure C.13).
@@ 5.3 Predicting Specific Relation Types
To apply the analysis conducted in the previous sections to real-world tasks, we focus on the relations between pairs of entity types that are of particular interest to practitioners (Gene-Gene, Drug-Gene and Drug-Disease). We investigate how they are represented in different biomedical KGs, and how these differences affect the predictive performance of KGE models. Statistics for the interactions of the considered entity types are given in Figure C.14, while model accuracy is plotted in Figure 10 (MRR) and Figure C.16 (Hits@10).
One basic way of characterising the performance of a KGE model is to consider if it can rank entities of the correct type above those of other types when predicting a specific relation (for instance, gene entities when predicting drug-gene interactions) – a task we refer to as demixing. As displayed in Figure C.15, the trained models tend to consistently predict entities of the correct type, albeit with different levels of accuracy. Interestingly, DistMult often shows worse demixing capabilities than other scoring functions. As a consequence of the generally strong demixing performance, the size of the potential set of tails for a given interaction needs to be taken into consideration when comparing model accuracy across different datasets and relation types (the smaller the set of candidates, the higher the expected MRR achieved by random ranking).
Gene-Gene. As shown in Figure 14(a), these interactions are characterized by a large number of potential tails and a large head out-degree of same relation type (with the exception of PharmKG for the latter), making predictions hard. However, a high predictive accuracy is observed on PharmKG and OpenBioLink: this is explained by the fact that, for these datasets only (Figure 14(b)), most triples are symmetric (and a significant portion also have inverse/inference edges), with the counterpart likely to have been seen during training. This also explains the relative ordering of scoring functions for OpenBioLink and PharmKG, with DistMult being the best and TransE the worst. Interestingly, ConvE also struggles on these highly symmetric datasets.
Drug-Gene. In Hetionet, PrimeKG and OpenBioLink these, interactions are easier to predict for the models than gene-gene interactions (Figure 10). This is likely due to an overall smaller number of potential tails seen during training and, in the case of PrimeKG, a markedly smaller head out-degree. PharmKG and PharMeBINet do not satisfy either of these properties, and indeed we observe no improvements in MRR/Hits@10. The strong predictive power observed for OpenBioLink can be linked to the presence of inverse edges for the vast majority of these triples.
@@ 5.4 Case Study: Effect of Additional Training Data
We consider eight relation types where we can find a significant number of common triples between the two datasets (statistics are given in Table C.1). We extract 10% of the shared triples of the considered relation type as test set and use all other edges in each graph for training. This requires retraining each KGE model separately for each investigated relation type. Due to the higher computational complexity of ConvE and its resemblance of the other investigated scoring functions throughout all results, we restrict this investigation to the four shallow embedding models. In the case of Hetionet, in addition to the embedding size maximizing memory utilization, we also train models with the same embedding size chosen for PharMeBINet, which is generally strictly smaller (Hetionet_max and Hetionet_same in Figure 11; hyperparameters in Table B.1). To further ensure a fair comparison, at test time we restrict predictions to a custom set of candidate tails that is the same for the two datasets, namely the set of entities appearing as tails in shared triples of the given relation type.
@@
We only report on MRR, as (similarly to previous sections) Hits@10 shows analogous patterns. As displayed in Figure 11 and Figure C.17, across all tested relation types there is no indication that the KGE models are able to benefit from the additional data seen when training on the larger PharMeBINet graph. On the contrary, models trained on Hetionet consistently perform better, even when using the same embedding size. Interestingly, the gap in MRR varies strongly across relation types and also across scoring functions, with DistMult showing generally little difference when comparing models trained with the same embedding size, while distance-based scoring functions experience worse degradation. From Table C.1 we notice that the relation types where the MRR gap is more significant tend to have a larger overall head out-degree (and more relation types coming out from head nodes) for triples in PharMeBINet, compared to Hetionet. We hypothesise that the fact that these nodes are used as head entities for more relation types and triples, many of which are likely not relevant for the specific task at hand, negatively impacts the quality of embeddings. On the other hand, Compound-downregulates-Gene and Compound-causes-Side Effect exemplify relation types where PharMeBINet contains far more triples in addition to the ones in Hetionet. Even in this case, where we could expect the additional training data in the larger dataset to be strictly relevant to the prediction task, all scoring functions show markedly degraded performance when trained on PharMeBINet.
@@ 6 Conclusions
This paper analyses the topological properties of widely used biomedical KGs and compares the corresponding predictive performance of different KGE models, focussing in particular on link-prediction tasks relevant to practitioners. Deviating from previous studies, we justify the need to look at properties of individual edges when trying to explain results, as pooling at the relation level introduces too much noise to detect any patterns. By going beyond the coarse-level one/many binary classification typically used for edge cardinality, we find that considering the actual degrees of head and tail nodes gives a stronger predictor explaining model accuracy. Interestingly, similar interpretations apply to both biomedical KGs and the general-domain KG. Edge topological patterns also impact predictive accuracy, especially when entity degrees are small. For such patterns, we observe an improved accuracy when the counterpart edge (e.g., the reverse edge for symmetric triples) has been seen during training. This in turn raises the problem of inconsistent model validation schemes and topological property definitions when comparing results of previous studies on similar datasets. We address this problem by releasing all predictions from our experiments, together with a new toolbox for KG topological analysis. Finally, by performing a case study comparing predictions on common sets of edges shared by different KGs, we show that training on larger graphs, encoding more biomedical data, can unintuitively harm predictive performance. This evidence should encourage a wider discussion on the guiding principles to adopt when constructing the training KG for biomedical tasks – a crucial problem for real-world practitioners, but scarcely investigated in the literature.
@@ Appendix A Dataset and Model Properties
Table A.2: Scoring functions and their ability to model four fundamental relation properties: S = Symmetry; INF = Inference; INV = Inversion; C = Composition. For RotatE we assume $d$ even and denote by $\mathbb{C}^{\frac{d}{2}}$ the vector space $\mathbb{R}^{d}=(\mathbb{R}\oplus i\mathbb{R})^{\frac{d}{2}}$ with the structure of $\mathbb{R}$-algebra induced by the product of complex numbers. $p\in\left\{1,2\right\}$; $\circ$ denotes the Hadamard product, w*w is the application of a convolutional filter $w$.
@@
TransE $\left\lVert{\bm{h}}+{\bm{r}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{r}},{\bm{t}}\in\mathbb{R}^{d}$
RotatE $\left\lVert{\bm{h}}\circ e^{i{\bm{r}}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{t}}\in\mathbb{C}^{\frac{d}{2}},{\bm{r}}\in\mathbb{R}^{\frac{d}{2}}$
@@
TripleRE $\left\lVert{\bm{h}}\circ\bm{r_{h}}+\bm{r_{m}}-{\bm{t}}\circ\bm{r_{t}}\right\rVert$ ${\bm{h}},\bm{r_{h}},\bm{r_{m}},{\bm{t}},\bm{r_{t}}\in\mathbb{R}^{d}$
ConvE $\langle\textrm{MLP}([{\bm{h}},{\bm{r}}]*w),{\bm{t}}\rangle$ ${\bm{h}},{\bm{r}},{\bm{t}}\in\mathbb{R}^{d},w\in\mathbb{R}^{3\times 3}$
@@ Appendix B Details on Experimental Setup and Hyperparameter Selection
All datasets were randomly split into training, validation and test set (8080% / 1010% / 1010%; in the case of PharMeBINet, 99.399.3% / 0.350.35% / 0.350.35% to mitigate the increased inference cost on the larger dataset). To ensure comparability across KGs, this random split was used even if pre-defined training, validation and test sets were provided with a dataset. We adopted log-sigmoid loss with negative adversarial sampling (Sun et al., 2019) and margin $12.0$, and the Adam optimiser (Kingma & Ba, 2015) for updating parameters. During training we always used negative sample sharing (Cattaneo et al., 2022). All experiments were performed on Graphcore IPUs using the BESS framework [^3] https://github.com/graphcore-research/bess-kge (Cattaneo et al., 2022). A fixed batch size of $768$ triples per device ($192$ for PharMeBINet) was adopted, while the embedding size for entities and relations was chosen for each KG and each scoring function independently to maximise the memory utilisation of a Bow-2000 IPU machine with 4 IPU processors (in the case of PharMeBINet, a Bow Pod16 with 16 IPUs). This is to ensure a fair comparison between scoring functions with different memory costs. For some scoring functions, especially DistMult, the memory footprint is typically dominated by the model parameters, allowing a larger hidden size for smaller KGs. For other scoring functions, especially TripleRE and ConvE, memory is typically dominated by activations, resulting in a similar hidden size for differently sized KGs. The learning rate, the norm used by the scoring function ($\mathrm{L}1$ or $\mathrm{L}2$) and the number of negative samples were determined by a hyperparameter sweep, based on the validation MRR (Table B.1). The full configurations can be found in the training scripts, released with the code [^4] https://github.com/graphcore-research/kg-topology-toolbox/tree/main/the_role_of_graph_topology_paper/train.
@@
Hetionet DistMult $2048$ $0.0003$ - $16$
RotatE $512$ $0.001$ $\mathrm{L}2$ $16$
TransE $1024$ $0.0001$ $\mathrm{L}1$ $16$
TripleRE $384$ $0.0001$ $\mathrm{L}1$ $16$
ConvE $676$ $0.001$ - $16$
Hetionet_same DistMult $300$ $0.0003$ - $16$
RotatE $128$ $0.003$ $\mathrm{L}2$ $16$
TransE $256$ $0.0003$ $\mathrm{L}1$ $16$
TripleRE $256$ $0.0001$ $\mathrm{L}1$ $16$
OpenBioLink DistMult $768$ $0.0003$ - $16$
RotatE $256$ $0.003$ $\mathrm{L}2$ $16$
TransE $512$ $0.0001$ $\mathrm{L}1$ $16$
TripleRE $256$ $0.001$ $\mathrm{L}2$ $16$
ConvE $529$ $0.001$ - $16$
PharMeBINet DistMult $300$ $0.003$ - $16$
RotatE $128$ $0.001$ $\mathrm{L}2$ $16$
TransE $256$ $0.00003$ $\mathrm{L}1$ $16$
TripleRE $256$ $0.0001$ $\mathrm{L}2$ $16$
ConvE $256$ $0.0001$ - $16$
PharmKG DistMult $768$ $0.003$ - $16$
RotatE $384$ $0.003$ $\mathrm{L}2$ $16$
TransE $768$ $0.0001$ $\mathrm{L}1$ $16$
TripleRE $384$ $0.0003$ $\mathrm{L}1$ $16$
ConvE $529$ $0.001$ - $16$
PrimeKG DistMult $1024$ $0.0003$ - $16$
RotatE $384$ $0.001$ $\mathrm{L}2$ $16$
TransE $768$ $0.0001$ $\mathrm{L}1$ $16$
TripleRE $256$ $0.0001$ $\mathrm{L}1$ $16$
ConvE $576$ $0.0003$ - $16$
FB15k-237 DistMult $4096$ $0.001$ - $16$
RotatE $1024$ $0.003$ $\mathrm{L}2$ $16$
TransE $2048$ $0.0001$ $\mathrm{L}1$ $16$
TripleRE $256$ $0.001$ $\mathrm{L}1$ $16$
ConvE $676$ $0.001$ - $16$